two rectangles are similar. the first is 4 in wide and 15 in long. the second is 9 in wide. find the length of the second rectangle.
Answer by oberobic(2304)

Since they are similar, you can use ratios to find the missing value.
.
4/15 = 9/x
.
Cross multiply.
4x = 135
.
Divide both sides by 4.
x=33.75
.
You can check by computing the ratios.
4/15 = .266667
.
9/33.75 = .266667
.
OK